If the cost of gas is given in the table below over these years. Calculate the inflation rate over the years listed in the table below?
Year Cost
1983 3.84
1984 3.56
1985 3.35
1986 3.78
Answer #1

The formula for yearly inflation rate is = (Cost of current year/Cost of previous year-1)*100

The inflation will be in percentage.

Year Cost Inflation Formula
1983 3.84
1984 3.56 -7.29 =(3.56/3.84-1)*100
1985 3.35 -5.9 =(3.35/3.56-1)*100
1986 3.78 12.84 =(3.78/3.35-1)*100